OpenStudy (anonymous):

Find the coordinates of the vertex for the parabola whose equation is f(x) = (x + 3)^2 + 3 (-3,3) (-3,-3) (0,3) (3,0)

OpenStudy (anonymous):

The vertex for this particular parabola is its low-point: Where x + 3 is a minimum or at x = -3 So, (-3, 3)

OpenStudy (anonymous):

This is in vertex form: \[ a ( x - h )^{2}+ k\] Where (h,k) is the vertex

OpenStudy (anonymous):

Example: \[a ( x - h) ^{2}+k\] \[3 ( x+ 5)^{2}+2\] The vertex would be -5, 2

OpenStudy (anonymous):

\[f(x) = (x + 3)^2 + 3 \] The easy thing that you should notice is that f(x)=y. All you have to do is replace f(x) with y and then you move the "+3" to the Left Hand Side(LHS). \[(x+3)^2= (y-3)\] You should know the vertex straight away because of this: \[(x-h)^2=(y-k)\] Where the point (h, k) is the vertex. Once you know that, you will know that (-3, 3) is the vertex in this question.